OpenStudy (anonymous):

Determine the distance from the origin with the points. Answer in mixed radical form (-6,-8) Please show me how to do this

OpenStudy (jdoe0001):

what are the coordinates of the "origin" anyway?

OpenStudy (anonymous):

Origin = (0,0)

OpenStudy (anonymous):

Make a triangle, then use the pythagorean theorem (A^2+B^2=C^2) to solve it. One length is 6. (0,0 to -6,0) The other side is 8 (-6,0 to -6,-8) and the last side is what we are trying to figure out. So now we have: 6^2+8^2=c^2 36+64=c^2 100=c^2 c=10 So the distance is 10. Sorry I don't have a picture to show you, it'd be easier to see that way.